# Efficient Online Reinforcement Learning Fine-Tuning Need Not Retain Offline Data

### v1

@@ 3.1 The Role of Offline Data at the Beginning of Fine-Tuning
Next, we wish to understand why recalibration leads to divergent Q-values. Figure 5 provides an intuition for
@@
such Q-value underestimation. Consider the very first batch of online rollouts collected from the environment. The target value computation for estimating TD-error on online state-action pairs will query the offline Q-function, $Q_{\theta}^{\mathrm{pre}}$, on out-of-distribution state-action pairs. Due to conservative offline RL pre-training, learned Q-values at these out-of-distribution state-action pairs are expected to be small. Using such pessimistic values for TD targets in the Bellman backup will, in turn, propagate these underestimation errors onto the new online state-action Figure 3(d) corroborates effective propagation of these TD targets. In addition, if one continues to run pessimistic RL algorithms during fine-tuning (e.g., CQL or CalQL), the conservative regularizer continues to minimize out-of-distribution Q-values: with few on-policy rollouts, Q-values for unseen actions keep getting smaller.
@@ 3.2 The Adverse Impact of Offline Data on Asymptotic Performance
Takeaway 3: Retaining offline data hurts asymptotic performance While retaining offline data appears to be crucial for preventing forgetting at the beginning of fine-tuning for current fine-tuning methods, continuing to make updates on this offline data with an (pesimistic) offline RL algorithm negatively impacts asymptotic performance and efficiency.
@@ 4 WSRL: Fast Fine-Tuning Without Offline Data Retention
WSRL algorithmWSRL is an off-policy actor-critic algorithm (Algorithm 1). It initializes the value function and policy with the pre-trained Q-function $Q_{\theta}^{\mathrm{pre}}$ and policy $\pi_{\psi}^{\mathrm{pre}}$ could come from any offline RL algorithm. Then, WSRL uses the first $K$ online steps to collect a few rollouts using the frozen offline RL policy to simulate the retention of offline data. We refer to this phase as the “warmup” phase. After warmup data collection, WSRL trains both the value and policy using standard temporal-difference (TD) updates and policy gradient. For fine-tuning, we fine-tune with a high updates-to-data (UTD) ratio [12, 5] and follow other best practices. To combat issues such as overestimation [19]in the high UTD regime, we use an ensemble of Q functions [5] and layer normalization [20] in both the actor and the critic.
@@ 5 Experimental Evaluation
The goal of our experiments is to study how well WSRL is able to fine-tune online without offline data retention. We also ablate the design decisions in WSRL to understand the reasons behind efficacy of WSRL. Concretely, we study the following research questions: (1) Can WSRL enable efficient finetuning in the no-retention setting?; (2) How does WSRL compare with methods that do retain offline data?; (3) How critical is the warmup phase in WSRL?; (4) How important is it to use online RL algorithm for online fine-tuning?, and (5) How important is it to pretrain the policy, value function, or both?
@@ 5.6 How important is it to initialize the policy, value function, and both?
Importance of policy initializaiton. At the start of online fine-tuning, WSRL initializes the policy to the pre-trained policy. Since the pre-trained policy is already capable of meaningfully acting in the environment, it speeds up online learning. In Figure 13, we compare WSRL’s performance with and without “policy initialization”, and find that initializing with the pre-trained policy is crucial for fast fine-tuning.
@@ Appendix C Experimental Setup
(1) The Antmaze tasks from D4RL [11] are a class of long-horizon navigation tasks that require controlling an $8$-DOF Ant robot to reach a goal with a sparse reward. The agent has to learn to “stitch” experiences together from a suboptimal dataset. In addition to the original mazes from D4RL, we include Antmaze-Ultra [24], a larger and more challenging maze. We only include three of the hardest Antmaze environments in Section 5, and provide results on all eight Antmazes in Appendix A. (2) The Kitchen environment is a long-horizon manipulation task to control a 9-DoF Franka robot arm to perfrom $4$ sequential subtasks in a simulated kitchen. (3) The Adroit environments are a suite of dexterous manipulation tasks to control a $28$-DoF five-fingered hand to manipulate a pen to desired position, open a door, and relocating a ball to desired position. The agent observes a binary reward when it succeeds. Each data has an offline dataset that provides a narrow offline dataset of $25$ human demonstrations and additonal trajectories collected by a behavior cloning policy. (4) The Mujoco Locomotion environments in D4RL are dense reward settings where agents learn to control robotic joints to perform various locomotion tasks.
@@ Appendix D Ablation Studies on Warmup Phase
Impact of different warmup types. One natural question arises: why does the simple approach of warming up the replay buffer significantly boost performance during fine-tuning? One hypothesis is that seeding the replay buffer with some data helps prevent early overfitting, as much work has found in online RL [34]. To test this hypothesis, we plot in Figure 16 in the fine-tuning performance of initializing with random actions, and compare it to initializing with pre-trained policy actions as well as not initializing the buffer atl all. It is clear that seeding the buffer with random actions significantly underperform the warmup approach, and in fact does not even provide much benefit as compared to not seeding the buffer at all. This suggests that the reason warmup phase helps is not because of preventing overfitting.
@@ Appendix E Unlearning and Recovery at the Start of Fine-tuning: An Analysis
Figure 20: Q-fucntion KL divergence $D_{KL}(softmax(Q_{\theta}^{\mathrm{pre}})||softmax(Q_{\theta}))$ between the pre-trained Q-function and the fine-tuned Q-function at the first $50k$ steps of fine-tuning. We evaluate the Q-functions by sampling states from the offline dataset distribution (Top) and the online buffer distribution (Bottom), and we sample actions by running $\pi_{\psi}^{\mathrm{pre}}$ and $\pi_{\psi}$ on the sampled states. WSRL is not plotted during the first $5,000$ steps of warmup. CQL and CalQL do not retain offline data.
@@ Appendix F Does Freezing the Policy at the Start of Fine-Tuning Help?
Figure 21: Freezing the policy for $N\in\{10k,30k\}$ steps at the onstart of fine-tuning doesn’t prevent the performance dip. In the plot, we show policy performance of WSRL vs. WSRL with initial policy freeze across six environments. Step 0 is the start of online fine-tuning.
@@ Appendix I Implementation Details
IQL Antmaze environments use expectile $0.9$, and all other environments use expetile $0.7$. For the inverse temperature in the actor, we use $10$ in Antmaze, $3$ in Mujoco locomotion, and $0.5$ for Kitchen and Adroit. Unless otherwise specified, IQL uses two Q functions and takes the mininum over them to estimate the Q value.
CQL, CalQL In Antmazes, we use the dual version of the CQL objective and set gap to $0.8$. In Mujoco locomotion and Kitchen environments, we set CQL regularizer weight $\alpha=5$. In Adroit environments, we set $\alpha=1$. CQL and CalQL use the same CQL regularizer for both pre-trainign and fine-tuning. In experiments that do retain offline data, each update batch samples $50\%$ from the offline data on Antmaze, Adroit, and Mujoco locomotion environments, and $25\%$ on Kitchen environments [33]. Unless otherwise specified, CQL/CalQL uses two Q functions and takes the mininum over them to estimate the Q-value.
RLPD / SAC(fast) We use an ensemble of $10$ Q-functions and a UTD of $4$ with batch size $256$. Following [3], in antmaze environments, we predict the Q-value from the ensemble by randomly subsampling $1$ Q-function. In all other environments, we subsample $2$ Q-functions and take the minimum over them to estimate the Q-value.
JSRL Same hyperparameters as RLPD, where we improve JSRL’s competitivenss with a Q-ensemble and UTD. For each online interaction episode, JSRL decides whether to roll in the pre-trained frozen policy or roll out the fine-tuned policy with probability $0.5$. This probability decreases linearly to 00 over the first $100k$ fine-tuning steps. In episode where the JSRL decides the roll in the pre-trained frozen policy, the number of step it rolls in follows a geometric distribution with $\gamma=0.99$, after which the fine-tuned policy is rolled out until the end of the episode.
SO2 Same hyperparameters as RLPD. Different from the original paper, we use $10$ Q-ensembles to make a fair comparison with WSRL and other baselines which have the same number of Q-ensembles. For the action noise, we use a standard normal with variance $0.3$, and we clip the action noise to be between $(-0.6,0.6)$.
@@
We investigate how WSRL performs with varying levels of expertise of the offline pre-trained policy. Specifically, we consider Kitchen-complete-v0 and Relocate-binary-v0, two especially hard tasks for offline RL where pre-training with CalQL leads to poor performance. In Recolate-binary-v0, CalQL completely fails and has pre-trained performance near 00; CQL and IQL also has pre-training performance 00, indicating that this task is inherently hard for offline RL agents. In Kitchen-complete-v0, CalQL ($15.47\%$) significantly underperforms IQL ($70.83\%$) despite our tunning efforts, which suggests there is some inherent limitation in CalQL learning a good Q-funciton in this domain. Not surprisingly, Figure 28 shows that WSRL also performs poorly: while WSRL can learn somewhat in Kitchen-complete-v0 with a non-zero initialization, it completely fails to learn in Adroit-binary-v0. This is expected because when pre-training fails, initializing with the pre-trained network may not bring any useful information gain, and may actually hurt fine-tuning by reducing the network’s plasticity [34], a known issue in online RL.

### v2

@@ 3.1 The Role of Offline Data at the Beginning of Fine-Tuning
Next, we wish to understand why recalibration leads to divergent Q-values. Figure 5 provides an intuition for such Q-value underestimation. Consider the very first batch of online rollouts collected from the environment.
@@
The target value computation for estimating TD-error on online state-action pairs will query the offline Q-function, $Q_{\theta}^{\mathrm{pre}}$, on out-of-distribution state-action pairs. Due to conservative offline RL pre-training (e.g. CQL or CalQL), learned Q-values at these out-of-distribution state-action pairs are expected to be small. Using such pessimistic values for TD targets in the Bellman backup will, in turn, propagate these underestimation errors onto the new online state-action Figure 3(d) corroborates effective propagation of these TD targets. In addition, if one continues to run pessimistic RL algorithms during fine-tuning (e.g., CQL or CalQL), the conservative regularizer continues to minimize out-of-distribution Q-values: with few on-policy rollouts, Q-values for unseen actions keep getting smaller.
@@ 3.2 The Adverse Impact of Offline Data on Asymptotic Performance
Takeaway 3: Retaining offline data hurts asymptotic performance While retaining offline data appears to be crucial for preventing forgetting at the beginning of fine-tuning for current fine-tuning methods, continuing to make updates on this offline data with an (pessimistic) offline RL algorithm negatively impacts asymptotic performance and efficiency.
@@ 4 WSRL: Fast Fine-Tuning Without Offline Data Retention
WSRL algorithm. WSRL is an off-policy actor-critic algorithm (Algorithm 1). It initializes the value function and policy with the pre-trained Q-function $Q_{\theta}^{\mathrm{pre}}$ and policy $\pi_{\psi}^{\mathrm{pre}}$ could come from any offline RL algorithm. Then, WSRL uses the first $K$ online steps to collect a few rollouts using the frozen offline RL policy to simulate the retention of offline data. We refer to this phase as the “warmup” phase. After warmup data collection, WSRL trains both the value and policy using standard temporal-difference (TD) updates and policy gradient. For fine-tuning, we fine-tune with a high updates-to-data (UTD) ratio [12, 5] and follow other best practices. To combat issues such as overestimation [19]in the high UTD regime, we use an ensemble of Q functions [5] and layer normalization [20] in both the actor and the critic.
@@ 5 Experimental Evaluation
The goal of our experiments is to study how well WSRL is able to fine-tune online without offline data retention. We also ablate the design decisions in WSRL to understand the reasons behind efficacy of WSRL. Concretely, we study the following research questions: (1) Can WSRL enable efficient fine-tuning in the no-retention setting?; (2) How does WSRL compare with methods that do retain offline data?; (3) How critical is the warmup phase in WSRL?; (4) How important is it to use online RL algorithm for online fine-tuning?, and (5) How important is it to pre-train the policy, value function, or both?
@@ 5.6 How important is it to initialize the policy, value function, and both?
Importance of policy initialization. At the start of online fine-tuning, WSRL initializes the policy to the pre-trained policy. Since the pre-trained policy is already capable of meaningfully acting in the environment, it speeds up online learning. In Figure 13, we compare WSRL’s performance with and without “policy initialization”, and find that initializing with the pre-trained policy is crucial for fast fine-tuning.
@@ Appendix C Experimental Setup
(1) The Antmaze tasks from D4RL [11] are a class of long-horizon navigation tasks that require controlling an $8$-DOF Ant robot to reach a goal with a sparse reward. The agent has to learn to “stitch” experiences together from a suboptimal dataset. In addition to the original mazes from D4RL, we include Antmaze-Ultra [24], a larger and more challenging maze. We only include three of the hardest Antmaze environments in Section 5, and provide results on all eight Antmazes in Appendix A. (2) The Kitchen environment is a long-horizon manipulation task to control a 9-DoF Franka robot arm to perfrom $4$ sequential subtasks in a simulated kitchen. (3) The Adroit environments are a suite of dexterous manipulation tasks to control a $28$-DoF five-fingered hand to manipulate a pen to desired position, open a door, and relocating a ball to desired position. The agent observes a binary reward when it succeeds. Each data has an offline dataset that provides a narrow offline dataset of $25$ human demonstrations and additional trajectories collected by a behavior cloning policy. (4) The Mujoco Locomotion environments in D4RL are dense reward settings where agents learn to control robotic joints to perform various locomotion tasks.
@@ Appendix D Ablation Studies on Warmup Phase
Impact of different warmup types. One natural question arises: why does the simple approach of warming up the replay buffer significantly boost performance during fine-tuning? One hypothesis is that seeding the replay buffer with some data helps prevent early overfitting, as much work has found in online RL [34]. To test this hypothesis, we plot in Figure 16 in the fine-tuning performance of initializing with random actions, and compare it to initializing with pre-trained policy actions as well as not initializing the buffer at all. It is clear that seeding the buffer with random actions significantly underperform the warmup approach, and in fact does not even provide much benefit as compared to not seeding the buffer at all. This suggests that the reason warmup phase helps is not because of preventing overfitting.
@@ Appendix E Unlearning and Recovery at the Start of Fine-tuning: An Analysis
Figure 20: Q-function KL divergence $D_{KL}(softmax(Q_{\theta}^{\mathrm{pre}})||softmax(Q_{\theta}))$ between the pre-trained Q-function and the fine-tuned Q-function at the first $50k$ steps of fine-tuning. We evaluate the Q-functions by sampling states from the offline dataset distribution (Top) and the online buffer distribution (Bottom), and we sample actions by running $\pi_{\psi}^{\mathrm{pre}}$ and $\pi_{\psi}$ on the sampled states. WSRL is not plotted during the first $5,000$ steps of warmup. CQL and CalQL do not retain offline data.
@@ Appendix F Does Freezing the Policy at the Start of Fine-Tuning Help?
Figure 21: Freezing the policy for $N\in\{10k,30k\}$ steps at the onset of fine-tuning doesn’t prevent the performance dip. In the plot, we show policy performance of WSRL vs. WSRL with initial policy freeze across six environments. Step 0 is the start of online fine-tuning.
@@ Appendix I Implementation Details
Code for WSRL is released at https://github.com/zhouzypaul/wsrl.
@@
IQL. Antmaze environments use expectile $0.9$, and all other environments use expectile $0.7$. For the inverse temperature in the actor, we use $10$ in Antmaze, $3$ in Mujoco locomotion, and $0.5$ for Kitchen and Adroit. Unless otherwise specified, IQL uses two Q functions and takes the minimum over them to estimate the Q value.
CQL / CalQL. In Antmazes, we use the dual version of the CQL objective and set gap to $0.8$. In Mujoco locomotion and Kitchen environments, we set CQL regularizer weight $\alpha=5$. In Adroit environments, we set $\alpha=1$. CQL and CalQL use the same CQL regularizer for both pre-trainign and fine-tuning. In experiments that do retain offline data, each update batch samples $50\%$ from the offline data on Antmaze, Adroit, and Mujoco locomotion environments, and $25\%$ on Kitchen environments [33]. Unless otherwise specified, CQL/CalQL uses two Q functions and takes the minimum over them to estimate the Q-value.
RLPD / SAC(fast). We use an ensemble of $10$ Q-functions and a UTD of $4$ with batch size $256$. Following [3], in antmaze environments, we predict the Q-value from the ensemble by randomly subsampling $1$ Q-function. This is needed for more optimism online in antmaze environments. In all other environments, we subsample $2$ Q-functions and take the minimum over them to estimate the Q-value.
JSRL. Same hyperparameters as RLPD, where we improve JSRL’s competitiveness with a Q-ensemble and UTD. For each online interaction episode, JSRL decides whether to roll in the pre-trained frozen policy or roll out the fine-tuned policy with probability $0.5$. This probability decreases linearly to 00 over the first $100k$ fine-tuning steps. In episode where the JSRL decides the roll in the pre-trained frozen policy, the number of step it rolls in follows a geometric distribution with $\gamma=0.99$, after which the fine-tuned policy is rolled out until the end of the episode.
SO2. Same hyperparameters as RLPD. Different from the original paper, we use $10$ Q-ensembles to make a fair comparison with WSRL and other baselines which have the same number of Q-ensembles. For the action noise, we use a standard normal with variance $0.3$, and we clip the action noise to be between $(-0.6,0.6)$.
@@
We investigate how WSRL performs with varying levels of expertise of the offline pre-trained policy. Specifically, we consider Kitchen-complete-v0 and Relocate-binary-v0, two especially hard tasks for offline RL where pre-training with CalQL leads to poor performance. In Recolate-binary-v0, CalQL completely fails and has pre-trained performance near 00; CQL and IQL also has pre-training performance 00, indicating that this task is inherently hard for offline RL agents. In Kitchen-complete-v0, CalQL ($15.47\%$) significantly underperforms IQL ($70.83\%$) despite our tunning efforts, which suggests there is some inherent limitation in CalQL learning a good Q-funciton in this domain. Not surprisingly, Figure 28 shows that WSRL also performs poorly: while WSRL can learn somewhat in Kitchen-complete-v0 with a non-zero initialization, it completely fails to learn in Adroit-binary-v0. This is expected because when pre-training fails, initializing with the pre-trained network may not bring any useful information gain, and may actually hurt fine-tuning by reducing the network’s plasticity [34], a known issue in online RL. In such situations, one may wish to resort to different methods (e.g. online RL) rather than fine-tuning from a bad initialization.